OpenStudy (anonymous):

Have to review trig. Can someone help me with: cos(x)-sin(x)=1

jimthompson5910 (jim_thompson5910):

cos(x)-sin(x)=1 (cos(x)-sin(x))^2 = 1^2 cos^2(x) - 2cos(x)*sin(x) + sin^2(x) = 1 (cos^2(x)+sin^2(x)) - 2cos(x)*sin(x) = 1 (1) - 2cos(x)*sin(x) = 1 1 - 2cos(x)*sin(x) = 1 I'll let you finish

OpenStudy (anonymous):

Find all solutions. I think I got it: cosx-sinx=1 (cosx-sinx)^2=(1)^2 cos^2x-sinxcosx-sincosx+sin^2x=1 cos^2x-2sinxcosx+sin^2x=1 (sin2x=2sinxcosx;double angle?) cos^2x-sin2x+sin^2x=1 (cos^2x => 1-sin^2x) 1-sin^2x-sin2x+sin^2x=1 -sin2x+1=1 sin2x=0 x=0,270

OpenStudy (anonymous):

Haha. I wish I would have looked at your reply before typing that out. -_- Thanks though!

jimthompson5910 (jim_thompson5910):

you are correct and yw, so it would be x = 0 + 360n where n is any integer or x = 270 + 360n where n is any integer

OpenStudy (tkhunny):

My favorite: \(\cos(x) - \sin(x) = \sqrt{2}\cos\left(x + \dfrac{\pi}{4}\right)\) This, of course, leads to the same solution.